OpenStudy (anonymous):

16a^2-25b^2 show work

OpenStudy (anonymous):

I assume you want it factored. Remember, a^2-b^2 = (a+b)(a-b) 16a^2-25b^2 = (4a - 5b)(4a+5b) (4a - 5b)(4a+5b)

OpenStudy (anonymous):

this difference of two squares = (4a+5b)(4a-5b)